OpenStudy (anonymous):

solve the system x+y=6 and x-y=2. What is the value of x? a. 6 b.2 c.4 d.0

OpenStudy (wasiqss):

x=4

Parth (parthkohli):

Adding the equations: 2x = 8 x = 4 y = 6

Parth (parthkohli):

y = 2 sorry

OpenStudy (callisto):

x+y=6 -(1) x-y=2 -(2) (1) +(2) 2x = 6+2 x=4 @wasiqss I know you're smart, but please show your solution so that the asker can learn :)

OpenStudy (anonymous):

thanks guys!:)

OpenStudy (wasiqss):

Sorry @Callisto i just woke up so dint feel like explaining all , but i will next time :)

Parth (parthkohli):

We know y = 2 as x - y = 2 ; 4 - y = 2 ; 4 = 2 + y ; 2 = y

OpenStudy (anonymous):

thanks parth Kohli!